OpenStudy (anonymous):

The function f(x) = 20(3)^x represents the growth of a fish population every year in a local lake. Jesse wants to manipulate the formula to an equivalent form that calculates every half-year, not every year. Which function is correct for Jesse's purposes?

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (mathstudent55):

\(f(x) = 20(3^x)\) Is this the correct function?

OpenStudy (anonymous):

no. It is f(x)=20(3)^x. Like a log function :)

jimthompson5910 (jim_thompson5910):

`Which function is correct for Jesse's purposes?` sounds like you're given a list of choices

OpenStudy (anonymous):

yes :)

OpenStudy (mathstudent55):

This function looks like it triples the number of fish each year.

OpenStudy (anonymous):

f(x)=20(3^2)^x/2 f(x)=20(3^1/2)^2x f(x)=20/2(3)^x (f(x)=40(2)^x these are the answer choices

OpenStudy (mathstudent55):

Let's look at the function from year 0 to year 3. Year 0: \(y = 20 \cdot 3^0 = 20 \cdot 1 = 20\) In year 0 there are 20 fish. Year 1: \(y = 20 \cdot 3^1 = 20 \cdot 3 = 60\) In year 1 there are 60 fish. Year 2: \(y = 20 \cdot 3^2 = 20 \cdot 9 = 180\) In year 2 there are 180 fish. Year 3: \(y = 20 \cdot 3^3 = 20 \cdot 27 = 540\) In year 3 there are 540 fish. As you can see, with each passing year, the population of fish triples.

OpenStudy (anonymous):

Yup! I see that! Continue please

OpenStudy (mathstudent55):

Now look at the given formulas. Since you are now calculating the growth half a year, you again start with a population of 20. After two half-years (which is 1 year), you must end up with 60. Try each choice with x = 2 (meaning 2 half-years = 1 year) to see which formula gives you 60.

OpenStudy (anonymous):

So would it be B?

OpenStudy (mathstudent55):

From what I wrote above, you know that f(0) = 20 f(1) = 60 f(2) = 180 f(3) = 540 Try each choice to see which one will give you those values, respectively, at f(0) f(2) f(4) f(6)

OpenStudy (mathstudent55):

Choice B. \(f(x)=20(3^{\frac{1}{2}})^{2x} = 20(3^x)\) \(f(0) = 20 \cdot 3^0 = 20\) \(f(2) = 20 \cdot 3^2 = 180\) It turns out that choice B s the same as the given function, so that is not it.

OpenStudy (mathstudent55):

Choice A. \(f(x)=20(3^2)^{\frac{x}{2}} = 20 \cdot 3^x\) Once again, just like choice B, choice A is also the same as the original function.

OpenStudy (anonymous):

Then that leaves choice C and D.

OpenStudy (anonymous):

So, C looks to be correct

OpenStudy (mathstudent55):

Choice C. \(f(x)=\dfrac{20}{2}(3)^x = 10(3^x) \) f(0) = 10 This cannot be correct because it gives us an initial value different from f(0) = 20

OpenStudy (mathstudent55):

Choice D. \(f(x)=40(2)^x\) \(f(0) = 20\) Once again, choice D also cannot be it because the initial value is wrong.

OpenStudy (anonymous):

so all of them are wrong??

OpenStudy (mathstudent55):

Since none of the choices you listed worked, I have a question for you. Notice that the functions in choices A and B were the same as the original expression. Are you sure you copied them correctly?

OpenStudy (anonymous):